OpenStudy (anonymous):

how do i find the square root of a negative number in terms of complex numbers. for example: sqrt(-64

OpenStudy (zzr0ck3r):

if \(a>0\) then \(\sqrt{-a}=i\sqrt{a}\) Example: \[\sqrt{-4}=i\sqrt{4}=i*2=2i\]

OpenStudy (anonymous):

so my answer would then be 8i?

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

thatnk!

OpenStudy (zzr0ck3r):

fo sure

OpenStudy (anonymous):

thanks*

OpenStudy (anonymous):

how do you type the square root symbol on the keyboard?